ax/2b - c/3d + x =1
Answer by tutorcecilia(2152)

ax/2b - c/3d + x =1 [Find the lowest common denominator: LCD= (3d)(2b) ]
.
(3d)(2b)(ax/2b) - (3d)(2b)(c/3d) + (3d)(2b)(x) =(3d)(2b)(1)
[Multiply each term by the LCD]
.
(3d)(ax) - (2b)(c) + (3d)(2b)(x) =(3d)(2b)(1)[Cancel wherever possible]
.
3adx-2bc+6bdx=6bd {I dont't think this can be simplified any further]